OpenStudy (anonymous):

pam received $100 for graduation and spends $6 of it a week . how many weeks can she do this and still have more than $25 left? this question hard it is inequalities

OpenStudy (anonymous):

100-6x>25 75-6x>0 75>6x 25>2x 12.5>x x=12

OpenStudy (bahrom7893):

Nice answer squee!

OpenStudy (anonymous):

i thouht that can it be 14 week

OpenStudy (anonymous):

Thanks. It's 12 because it have to be an integer, it isn't 13 because it has to be LESS than 12.5

OpenStudy (anonymous):

14*6=84 100-84=16 25>16

OpenStudy (anonymous):

how you to graph it

OpenStudy (anonymous):

I'm not really sure, so I'll let someone else answer instead to risking to give you a wrong answer I never was good with graphs.